all questions correctly in pen in the esamination booklet (10 pts cack) Hs pressure are present in a sample which occupies 450 ml at 32C, 720 mm 2) W hat is molar the molar mass ofagas with a density of 6.29 g/L at 1 atm, 25°C mixture of 14.2 g He and 36.7 g Ar, is placed in a 100 mL container at 290 a) what is the total pressure of the system? b) what is the mole fraction of each gas? e) what is the partial pressure of each gas? 4) How many grams of HgO need to react to produce 11.5 liters of O, at 750 mm He,3MPC by 5) An unknown gas was found to effuse 2.92 times slower than Hajp under identical conditions What is the molar mass of the unknown?


plz answer all questions
Answer #1

1.PV = nRT
   n = W/M
PV = WRT/M
P = 720/760   = 0.947atm
T = 32+273     = 305K
V = 0.45L
M = 44g/mole
W = PVM/RT
      = 0.947*0.45*44/0.0821*305   = 0.75g
2.PM = dRT
   d = 6.29g/L
   P   = 1atm
   T   = 25+273 = 298K
   M   = dRT/P
       = 6.29*0.0821*298/1 = 154g/mole
3. no of moles of He = 14.2/4 = 3.55moles
   no of moles of Ar = 36.7/40 = 0.9175 moles
   total no of moles of He,Ar(n)= 3.55+0.9175 = 4.4675moles
                              v = 0.1L
                               T = 290K
          PV =nRT
           P = nRT/V
              = 4.4675*0.0821*290/0.1    = 1063.66atm
    mole fraction of He = no of moles He/total no of moles
                         = 3.55/4.4675 = 0.794
     mole fraction of Ar = no of moles He/total no of moles
                         = 0.9175/4.4675 = 0.2053
    partial pressure of He = mole frarion of He * total pressure
                            = 0.794*1063.66 = 844.54atm
    partial pressure of Ar = mole frarion of Ar * total pressure
                            = 0.2053*1063.66 = 218.36atm